OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form. –2x – y = –7 2x – y = –7 2x – y = 11 –2x + y = –11

OpenStudy (anonymous):

Standard form is ax+by=c so where would your (x,y) and slope go?

OpenStudy (anonymous):

y-y0 = m(x-x0) y-(-1)=2(x-5) y+1=2x-10 therefore, 2x-y = 11